# LLM Critics Help Catch Bugs in Mathematics: Towards a Better Mathematical Verifier with Natural Language Feedback

### v3

@@ Abstract
Mathematical verfier achieves success in mathematical reasoning tasks by validating the correctness of solutions. However, existing verifiers are trained with binary classification labels, which are not informative enough for the model to accurately assess the solutions. To mitigate the aforementioned insufficiency of binary labels, we introduce step-wise natural language feedbacks as rationale labels (i.e., the correctness of the current step and the explanations). In this paper, we propose Math-Minos, a natural language feedback enhanced verifier by constructing automatically-generated training data and a two-stage training paradigm for effective training and efficient inference. Our experiments reveal that a small set (30k) of natural language feedbacks can significantly boost the performance of the verifier by the accuracy of 1.6% (86.6% $\rightarrow$ 88.2%) on GSM8K and 0.8% (37.8% $\rightarrow$ 38.6%) on MATH. We have released our code and data for further exploration.
@@ 1 Introduction
Figure 1: Illustration of Score Feedback and Natural Language Feedback. $q$ represents the mathematical questions $s_{1},s_{2},s_{3}$ represent the intermediate solutions. $a$ represents the final answer. 00 and $1$ represent the score feedbacks. Our work aims to mitigate the insufficiency of score feedbacks and enhance verifiers’ evaluation capabilities by introducing step-wise natural language feedbacks.
Large Language Models (LLMs) Bai et al. (2023); Touvron et al. (2023a, b); Jiang et al. (2023); OpenAI et al. (2024) have demonstrated strong capabilities in summarization Touvron et al. (2023b), coding Rozière et al. (2024), tool using Song et al. (2023) and dialogue Ouyang et al. (2022). However, mathematical reasoning remains a challenge for LLMs (Lightman et al., 2023; Huang et al., 2024). To tackle this problem, recent research has focused on using verifiers to validate the correctness of response generated by models (Wang et al., 2023b; Zhu et al., 2023; Li et al., 2023; Wang et al., 2024). An effective verifier can serve as 1) response reranker in the decoding Li et al. (2023); Yu et al. (2024a); Wang et al. (2024). 2) reward model in RLHF Shao et al. (2024); 3) data purifier that filters erroneous responses in the SFT Rafailov et al. (2023).
@@
In this work, we aim to enhance the verifier’s evaluation ability for mathematical solutions by introducing step-level natural language feedbacks as rationale labels (i.e., the correctness of the current step and the explanations). We propose MATH-Minos, a natural language feedback enhanced verifier as  section 3. By employing supervised fine tuning on only 30k training data with natural language feedbacks as rationale labels before binary classification training, we can effectively enhance the model’s evaluation capabilities. In the first stage, we create high-quality step-level natural language feedback data as  subsection 3.2. In order to address the challenge of accurate evaluation generation, we introduce Label-aware Natural Language Feedback Curation to simplify the task by introducing step-level binary classification labels to enhance GPT-4’s evaluation generation. The natural language feedbacks can provide in-depth reasons behind classification feedbacks, which is helpful for training the verifier. In the second stage, we introduce a two-stage training for MATH-Minos as subsection 3.3: firstly, we adopt the supervised fine tuning on rationale labels to effectively help improve the model’s evaluation capabilities, followed by standard ORM & PRM training on score feedbacks to achieve efficient inference with a single forward step.
The experiments in  section 4 demonstrate that infusing the model with evaluation capabilities via natural language feedback sas rationale labels is more efficient and effective than score feedbacks. We show that only 30k training data with natural language feedbacks can significantly boost the performance of mathematical verifiers. For Outcome Reward Model (ORM) setting, MATH-Minos improves the accuracy of MetaMath-Mistral (Yu et al., 2024b) by 1.6% (86.6% $\rightarrow$ 88.2%) on GSM8K and 0.7% on MATH % (37.6% $\rightarrow$ 38.3%) for ORM. For Process Reward Model (PRM) setting, MATH-Minos improves the accuracy by 0.7% (87.1% $\rightarrow$ 87.8%) on GSM8K and 0.8% (37.8% $\rightarrow$ 38.6%) on MATH.
@@
1. We are the first to conduct in-depth analyses on the reasons behind incorrect evaluations generated by verifiers, revealing the shortcomings of current verifier’s training paradigm and inspiring future research.
2. We propose and demonstrate that training verifiers with natural language feedbacks can complement the non-informative score feedbacks thus enhancing the model’s evaluation ability.
3. We propose MATH-Minos by proposing label-aware natural language feedback curation and two-stage training paradigm.
@@ 2 Related Works
Previous works focus on improving the mathematical reasoning ability of LLMs on three ways: (1) Pre-training: LLMs Azerbayev et al. (2023); OpenAI et al. (2024); Touvron et al. (2023a, b) are pre-trained on a large set of corpus related to mathematical questions with next-token prediction objective. (2) Supervised fine-tuning: Supervised fine-tuning can also improve the mathematical reasoning ability of LLMs by training LLMs with mathematical questions with detailed solutions  Yu et al. (2023); Luo et al. (2023); Wang et al. (2023a). (3) Inference:  Wei et al. (2022); Fu et al. (2022); Zhang et al. (2023); Bi et al. (2023) design prompting strategies to improve the reasoning ability of LLMs.
@@
Previous mathematical verifiers can be mainly categorized into two categories: Outcome Reward Model (ORM) gives a reward to the whole solution; Process Reward Model (PRM) gives a reward to each reasoning step of the solution. Previous works Li et al. (2023); Ying et al. (2024); Wang et al. (2024) use question-solution pair data with a binary label to train a ORM or a PRM, which is inefficient to help models understand the errors. In this work, we aim to leverage step-wise natural language feedback to enhance the training of traditional verifiers.
@@ 3.2 Label-aware Natural Language Feedback Curation
In this section, we introduce the label-aware natural language feedback curation of MATH-Minos as shown in Figure 2. Since the natural language feedback can be understood by both humans and large models, it is suitable to stimulate the evaluation capabilities of LLMs. Unlike the binary label, natural language feedback provides detailed explanations for right or wrong completions, which also brings complexity to data collection. The best way for generating the natural language feedback data is through manual annotation. Considering the costs associated with human annotation, we obtain natural language feedback by leveraging the capabilities of the most advanced LLM, GPT-4-turbo (OpenAI et al., 2024).
@@
To facilitate GPT-4 in generating higher quality data, we propose a label-aware prompting method, which simplify the evaluation task by introducing the binary classification label within the prompt. As illustrated in Figure 2, GPT-4’s task shifts from determining correctness and generating explanations to generating explanations based on the given label. Extensive analysis in Section 5 have also validate the effectiveness of our approach.
@@ 3.3 Two-Stage Training of MATH-Minos
After the first stage, the evaluation capability of the model is improved. However, natural language feedback cannot provide a score and thus can’t be used as a reward for further optimizations like PPO. Schulman et al. (2017) Additionally, when the model generates feedback, it produces a complete evaluation with rationales, making it significantly less efficient than using a classification-based verifier. Therefore, we further train the verifier with binary classification labels as Equation 1 and Equation 2.
@@ 4.1 Experiment Setup
Following Li et al. (2023) and Wang et al. (2024), we also explore the ensemble of self-consistency (majority voting) and the verifier. Specifically, we classify the results output by the model into different groups and calculate the cumulative reward for each group, which can be calculate as follows:
@@
For the training of the verifiers, to ensure comparability and convenience, we utilize the open-source MATH-shepherd dataset Wang et al. (2024) for both baseline reward models and MATH-Minos. We curate total 30K samples of natural language feedback using the data from phase-one of PRM800K Lightman et al. (2023) and the subset of MATH-Shepherd. Our main experiment conducts the verification on the test set of GSM8K and MATH. We use the MetaMATH-Mistral as the generator for the questions in the test set. In order to ensure the model has the ability of solving mathematical problem before learning to evaluate, we also use the MetaMATH-Mistral as the base model for MATH-Minos and all other reward models. For the training of natural language feedback, we use 30k training data generated as subsection 3.2 with learning rate of 5e-6 with total batch size of 256. For the training of score feedback, we use 440k training data (i.e., 30k data with the binary classification labels from the training data in the training of natural language feedbacks and 410k data sampled from MATH-Shepherd. For the training of baseline, we use the total 440k training data from MATH-Shepherd. For the training of ORM, we adopt the learning rate of 3e-6 with the batch size of 512. For PRM, the learning rate is 2e-6 with the batch size of 512.
@@ 4.2 Main Result
We present the performance of various methods in Table 2. Compared to traditional ORM, MATH-Minos (ORM) achieves an improvement of 1.1% in accuracy on the GSM8K and 0.7% in on the MATH. For PRM, MATH-Minos (PRM) achieves an accuracy improvement of 0.7% on GSM8K and 0.8% on the MATH. Ensembling with self-consistency and MATH-Minos, the MetaMATH-Mistral generator achieves optimal accuracy of 88.2% on GSM8K and 38.6% on MATH500.
Beyond the improvement of the performance, we find that MATH-Minos has a more pronounced effect in the setting of ORM. We believe this phenomenon could be attributed to the sparser supervision in ORM compared to PRM, implying that information-rich textual explanations can offer more substantial benefits to ORM.
@@ 5 Analysis
5.1 Error Distributions of the Math Solvers
To further illustrate the shortcomings of training the verifier solely relying on binary classification, we conduct an in-depth investigation into the errors produced by the generator at the step level. Specifically, we take the natural language feedback generated by GPT-4 as a reference and heuristically and categorized the causes of errors in responses into five distinct types: Unrelated: This indicates that the step is irrelevant and does not contribute towards deducing the final answer. Accumulation: This denotes that the step is incorrect due to a mistake in the preceding step, leading to subsequent errors. Calculation: This categorization is reserved for errors arising from incorrect calculations, which is one of the most common errors in mathematical reasoning. Logic: This applies to steps that are logically flawed in the context of solving the given problem. Other: This category encompasses steps that are erroneous for reasons not covered by the aforementioned categories.
@@ 5.1 Error Distributions of the Math Solvers
We use the same way as the label-aware prompting introduced in Section 2 to automatically analyze the cause of errors. We obtain the step-level labels from the subset of MATH-Shepherd (Wang et al., 2024), which contains 500 samples for both GSM8K and MATH. Given the question, solution and the natural language feedbacks, we employ GPT-4 for the classification of error causes. The experimental result is shown in Figure 4.
From our statistical analysis, it is evident that the model produces errors across all types. For the MATH dataset, given it higher difficulty level and the necessity for more steps, a greater total number of errors occur within the same number of samples of GSM8K. Furthermore, we discovere that the most common cause of errors in both datasets is accumulation, which is consistent with our intuition. In multi-step reasoning, a mistake in one step is likely to directly cause errors in all subsequent steps. Furthermore, we observe distinct patterns of errors in the GSM8K and MATH datasets. For the GSM8K dataset, the occurrences of calculation errors and logical errors were approximately the same. Instead, in the MATH dataset, logical errors significantly outnumber calculation errors. This also indirectly demonstrates that models are vulnerable in more complex reasoning tasks.
@@ 5.2 Meta-Evaluation and Convergence Curves
Figure 5: The convergence curve of vanilla ORM trained with score feedbacks (i.e., 440k quesiton-solution data with classification labels) and our Math-Minos with natural language feedbacks (i.e., 30k quesiton-solution data with rationale labels) on GSM8K.
To measure the verifier’s capabilities in a more convenient and direct manner instead of verification, a intuitive approach is to assess whether the verifier can accurately determine the correctness of the final answer. Without the influence of the generator, this method purely relies on the capability of the verifier. We construct a meta-evaluation set based on whether the final answer provided in the generator’s output is correct, serving as the ground truth label (despite the potential for false positives). By sampling several responses from the generator on the test set and deriving labels through rules, we create a meta-evaluation set for GSM8K containing 20,000 samples. We conduct tests on the meta-evaluation set using the checkpoints of each epoch after completing the model training and verification. The results of the meta-evaluation are presented in Figure 5.
@@
5.3 Influence of the Data Quality
@@
To validate the effectiveness of Label-Aware Natural Language Feedback Curation, we conduct a comparative experiment against directly prompting. We use the 30K direct GPT-4 evaluation which is the same number of MATH-Minos to compare. We use both meta-eval and verification to measure the capability of the verifier. The experimental result is shown in Table 3.
The experimental results indicate that directly prompting GPT does not significantly enhance the performance of the verifier. This is possibly due to the quality of the data shown in Table 1. The hallucinations produced by GPT-4 can further accumulate in the verifier, thus affecting the final performance of the model.

### v4

@@ Abstract
In recent progress, mathematical verifiers have achieved success in mathematical reasoning tasks by validating the correctness of solutions generated by policy models. However, existing verifiers are trained with binary classification labels, which are not informative enough for the model to accurately assess the solutions. To mitigate the aforementioned insufficiency of binary labels, we introduce step-wise natural language feedback as rationale labels, that is, the correctness of each step and the detailed explanations. In this paper, we propose Math-Minos, a natural language feedback-enhanced verifier by constructing automatically generated training data and a two-stage training paradigm for effective training and efficient inference. Our experiments reveal that a small set of natural language feedback can significantly boost the performance of the verifier in both verification and reinforcement learning. We have released the code and data for further exploration.
@@ 1 Introduction
Figure 1: Illustration of Score Feedback and Natural Language Feedback. $q$ represents the mathematical questions $s_{1},s_{2},s_{3}$ represent the intermediate solutions. $a$ represents the final answer. 00 and $1$ represent the score feedback. Our work aims to mitigate the insufficiency of score feedback and enhance verifiers’ evaluation capabilities by introducing step-wise natural language feedback.
Large Language Models (LLMs) Bai et al. (2023); Touvron et al. (2023a, b); Jiang et al. (2023); OpenAI et al. (2024) have demonstrated strong capabilities in summarization Touvron et al. (2023b), coding Rozière et al. (2024), tool using Song et al. (2023) and dialogue Ouyang et al. (2022). However, mathematical reasoning remains a challenge for LLMs Lightman et al. (2023); Huang et al. (2024). To tackle this problem, recent research has focused on using verifiers to validate the correctness of response generated by models (Wang et al., 2023b; Zhu et al., 2023; Li et al., 2023; Wang et al., 2024). An effective verifier can serve as 1) response reranker in the decoding Li et al. (2023); Yu et al. (2024a); Wang et al. (2024); 2) reward model in further post-training Shao et al. (2024); Wang et al. (2024); 3) data purifier that filters erroneous responses in the SFT Rafailov et al. (2023).
@@
In this work, we aim to enhance the verifier’s evaluation ability for mathematical solutions by introducing step-level natural language feedback as rationale labels (i.e., the correctness of the current step and the further explanations). We propose MATH-Minos, a natural language feedback enhanced verifier as shown in Section 3. By employing supervised fine-tuning on only 30k natural language feedback before binary classification training, we can effectively enhance the model’s evaluation capabilities. In the first stage, we create high-quality step-level natural language feedback data as Section 3.2. To improve the quality of natural language evaluation, we introduce Label-aware Natural Language Feedback Curation. This approach simplifies the task by incorporating step-level binary classification labels, enhancing the evaluation generation process for GPT-4. Natural language feedback can provide in-depth reasons behind classification feedback, which enhances the training of the verifier. In the second stage, we introduce a two-stage training for MATH-Minos as Section 3.3: Firstly, we adopt the supervised fine-tuning on curated natural language feedback to effectively help improve the model’s evaluation capabilities, followed by standard ORM & PRM training on score feedbacks to achieve efficient inference with a single forward step.
The experiments in Section 4 demonstrate that infusing the model with evaluation capabilities via natural language feedback is more efficient and effective than traditional score feedback. We show that only 30k training data with natural language feedback can significantly boost the performance of mathematical verifiers. For Outcome Reward Model (ORM) setting, MATH-Minos improves the accuracy of MetaMath-Mistral (Yu et al., 2024b) by 1.6% (86.6% $\rightarrow$ 88.2%) on GSM8K and 0.7% on MATH % (37.6% $\rightarrow$ 38.3%) for ORM. For Process Reward Model (PRM) setting, MATH-Minos improves the accuracy by 0.7% (87.1% $\rightarrow$ 87.8%) on GSM8K and 0.8% (37.8% $\rightarrow$ 38.6%) on MATH. For reinforcement learning, MATH-Minos serves as a better reward model with an accuracy improvement by 1.8% (81.8% $\rightarrow$ 83.6%) on GSM8K and 1.5% (31.3% $\rightarrow$ 32.8%) on MATH.
@@
1. We are the first study to conduct in-depth analyses of the reasons behind the evaluations generated by verifiers, revealing the shortcomings of the current verifier’s training paradigm and inspiring future research.
2. We propose MATH-Minos by proposing label-aware natural language feedback curation and a two-stage training paradigm, demonstrating that training verifiers with natural language feedback can complement the non-informative score feedback thus enhancing the model’s evaluation ability.
@@ 2 Related Works
Previous works focus on improving the mathematical reasoning ability of LLMs on three ways: (1) Pre-training: LLMs Azerbayev et al. (2023); OpenAI et al. (2024); Touvron et al. (2023a, b) are pre-trained on a large set of corpus related to mathematical questions with next-token prediction objective. (2) Supervised fine-tuning: Supervised fine-tuning can also improve the mathematical reasoning ability of LLMs by training LLMs with mathematical questions with detailed solutions  Yu et al. (2023b); Luo et al. (2023); Wang et al. (2023a). (3) Inference:  Wei et al. (2022); Fu et al. (2022); Zhang et al. (2023); Bi et al. (2023) design prompting strategies to improve the reasoning ability of LLMs.
@@
Previous mathematical verifiers can be mainly categorized into two categories: Outcome Reward Model (ORM) gives an evaluation score to the whole solution; Process Reward Model (PRM) gives an evaluation score to each intermediate step of the solution. Previous works Yu et al. (2023a); Ying et al. (2024); Wang et al. (2024) use question-solution pair data with a score to train a ORM or a PRM, which is inefficient to help models understand the errors. Therefore, in this work, we aim to train a verifiers with error types and detailed explanations about the errors.
@@ 3.2 Label-aware Natural Language Feedback Curation
In this section, we introduce the label-aware natural language feedback curation of MATH-Minos as shown in Figure 2. Since the natural language feedback can be understood by both humans and large models, it is suitable to stimulate the evaluation capabilities of LLMs. Unlike the binary label, natural language feedback provides detailed explanations for right or wrong completions, which also brings complexity to data collection. The best way for generating the natural language feedback data is through manual annotation. Considering the costs associated with human annotation, we obtain natural language feedback by leveraging the capabilities of the GPT-4-turbo (OpenAI et al., 2024).
@@
To facilitate GPT-4 in generating higher-quality data, we propose a label-aware prompting method, which simplifies the evaluation task by introducing the binary classification label within the prompt. As illustrated in Figure 2, GPT-4’s task shifts from determining correctness and generating explanations to generating explanations based on the given label. After obtaining automatically labeled data, we hire a group of graduate students to review the data, further ensuring its quality.
@@ 3.3 Two-Stage Training of MATH-Minos
After the first stage, the evaluation capability of the model is improved. However, natural language feedback cannot provide a score and thus can’t be used as a reward for further optimizations like PPO Schulman et al. (2017). Additionally, when the model generates feedback, it produces a complete evaluation with rationales, making it significantly less efficient than using a classification-based verifier. Therefore, we further train the verifier with binary classification labels as Equation 1 and Equation 2.
@@ 4.1 Experiment Setup
Following Li et al. (2023) and Wang et al. (2024), we also explore the ensemble of self-consistency (majority voting) and the verifier. Specifically, we classify the results output by the model into different groups and calculate the cumulative reward for each group, which can be calculated as follows:
@@
Reward Model
Following Wang et al. (2024), to further examine the performance of the verifier, we evaluate it as the reward model for PPO (Schulman et al., 2017) during the post-training phase. An effective reward model can provide reliable feedback signals, thereby enhancing the final performance of PPO.
@@
For the training of the verifiers, to ensure comparability and convenience, we utilize the open-source MATH-shepherd dataset Wang et al. (2024) for both baseline reward models and MATH-Minos. We curate a total of 30K samples of natural language feedback using the data from phase-one of PRM800K Lightman et al. (2023) and the subset of MATH-Shepherd. Our main experiment conducts the verification on the test set of GSM8K and MATH and reinforcement learning (PPO) on the training set of GSM8K and MATH. We follow Wang et al. (2024) with their detailed experimental setup, as shown in Appendix D.
Model GSM8K MATH500
Mistral-7B: MetaMATH 77.9 28.6
+ ORM-PPO 81.8 31.3
+ MATH-Minos(ORM)-PPO 83.6 32.8
Table 3: Experimental result of the MATH-Minos in PPO post-training setting. Our MATH-Minos significantly outperforms the vanilla ORM on PPO training.
@@ 4.2 Main Result
We present the performance of various methods in Table 2 and Table 3.
For the task of verification, MATH-Minos (ORM) achieves an improvement of 1.1% in accuracy on the GSM8K and 0.7% in on the MATH compared to traditional ORM. For PRM, MATH-Minos (PRM) achieves an accuracy improvement of 0.7% on GSM8K and 0.8% on the MATH. Ensembling with self-consistency and MATH-Minos, the MetaMATH-Mistral generator achieves optimal accuracy of 88.2% on GSM8K and 38.6% on MATH500. Beyond the improvement of the performance, we find that MATH-Minos has a more pronounced effect in the setting of ORM. We believe this phenomenon could be attributed to the sparser supervision in ORM compared to PRM, implying that information-rich textual explanations can offer more substantial benefits to ORM.
For the task of reward model on reinforcement learning, MATH-Minos (ORM) achieves an improvement of 1.8% in accuracy on the GSM8K and 1.5% in on the MATH compared to traditional ORM, demonstrating the robust effectiveness of our method across various task settings.
@@ 5 Analysis
5.1 Error Distribution of Math Solvers
To further illustrate the shortcomings of training the verifier solely relying on binary classification, we conduct an in-depth investigation into the errors produced by the generator at the step level. Specifically, we take the natural language feedback generated by GPT-4 as a reference and heuristically categorize the causes of errors in responses into five distinct types: Unrelated Unrelated: This indicates that the step is irrelevant and does not contribute towards deducing the final answer. Accumulation: This denotes that the step is incorrect due to a mistake in the preceding step, leading to subsequent errors. Calculation: This categorization is reserved for errors arising from incorrect calculations, which is one of the most common errors in mathematical reasoning. Logic: This applies to steps that are logically flawed in the context of solving the given problem. Other: This category encompasses steps that are erroneous for reasons not covered by the aforementioned categories.
@@ 5.1 Error Distribution of Math Solvers
We use the same way as the label-aware prompting introduced in Section 2 to automatically analyze the cause of errors. We obtain the step-level labels from the subset of MATH-Shepherd (Wang et al., 2024), which contains 2500 samples for both GSM8K and MATH. Given the question, solution and the natural language feedback, we employ GPT-4 for the classification of error causes. The experimental result is shown in Figure 4.
From our statistical analysis, it is evident that the model produces errors across all types. For the MATH dataset, given its higher difficulty level and the necessity for more steps, a greater total number of errors occur within the same number of samples of GSM8K. Furthermore, we discover that the most common cause of errors in both datasets is accumulation, which is consistent with our intuition. In multi-step reasoning, a mistake in one step is likely to directly cause errors in all subsequent steps. Furthermore, we observe distinct patterns of errors in the GSM8K and MATH datasets. For the GSM8K dataset, the occurrences of calculation errors and logical errors were approximately the same. Instead, in the MATH dataset, logical errors significantly outnumber calculation errors. This also indirectly demonstrates that models are vulnerable to more complex reasoning tasks.
@@ 5.2 Meta-Evaluation and Convergence Curves
To measure the verifier’s capabilities in a more convenient and direct manner instead of verification, an intuitive approach is to assess whether the verifier can accurately determine the correctness of the final answer. Without the influence of the generator, this method purely relies on the capability of the verifier. We construct a meta-evaluation set based on whether the final answer provided in the generator’s output is correct, serving as the ground truth label. By sampling several responses from the generator on the test set and deriving labels through rules, we create a meta-evaluation set for GSM8K containing 20,000 samples. We conduct tests on the meta-evaluation set using the checkpoints of each epoch after completing the model training and verification. The results of the meta-evaluation are presented in Figure 6. Additionally, we construct the step-level meta-evaluation dataset using the same data in Section 5.1 to provide insights into where our method’s gains are.
Figure 5: The convergence curve of vanilla ORM and our Math-Minos with natural language feedbacks on meta-evaluation set.
Figure 6: The detailed improvement of the step-level model prediction accuracy between vanilla PRM and Math-Minos.
@@
Furthermore, in the step-level meta-evaluation as Figure 6, we can observe that our method significantly aids in identifying all types of errors, with the most notable improvements seen in logical errors and unrelated contents. This aligns well with the strengths of natural language feedback, which provides a more detailed analysis of the underlying reasons compared to score-based feedback. Such insights help enhance the logical reasoning capabilities of the reward model and effectively identify redundant information.
5.3 Influence of the Feedback Quality
@@
GSM8K GSM8K MATH
Vanilla ORM 86.2 35.9
MATH-Minos (GPT-4) 87.3 37.4
MATH-Minos (Qwen2-72b) 87.5 36.8
MATH-Minos (Mistral-7B) 85.4 34.5
Table 5: Experimental result of the ORM, Math-Minos with label-aware natural language feedback curation from different models. "Qwen2-72b" denotes we use Qwen2-72b-instruct as the natural language feedback generator. "Mistral-7B" denotes we use Mistral-7b-Instruct-v0.3 as the natural language feedback generator.
In this section, we conduct an extensive analysis of the impact on the data quality of natural language feedback. One baseline is direct prompting of GPT-4-turbo. We use the 30K direct GPT-4 evaluation which is the same number of MATH-Minos to compare. We use both meta-eval and verification to measure the capability of the verifier. Additionally, to enhance the scalability and accessibility of our method, we explore the use of different base models to generate natural language feedback. Specifically, we utilize the open-source model Qwen2-72b-Instruct as a substitute for GPT-4-turbo. Furthermore, we take a more radical approach by directly employing the identical base model, Mistral-7B-Instruct-v0.3, which shares the same base model of our policy model MetaMATH: Mistral-7B. This allows us to comprehensively investigate the impact of the natural language feedback generated by these models on the verifier’s performance.The experimental result is shown in Table 4 and Table 5.
The experimental results indicate that directly prompting GPT does not significantly enhance the performance of the verifier. This is possibly due to the poor quality of the data shown in Table 1. From the experimental results in Table 5, we observed that Qwen2-72b-Instruct can produce high-quality natural language feedback, leading to a similar enhancement of GPT-4-turbo in the verifier’s performance, thereby demonstrating the scalability and accessibility of our approach. However, we found that Mistral-7B-Instruct-v0.3, which shares the same base model as our verifier, fails to improve the model’s performance effectively. We believe this may be due to the model’s smaller size, which hinders its ability to generate accurate natural language feedback, ultimately resulting in a decline in the verifier’s performance.